Two numbers are such that their difference, their sum and their product are to be one another as 1:7:24. What is the product of the two numbers?
Answer by checkley77(12844)

1:7:24
x-y=1
x+y=7
xy=24
------------
8-6=2:1
8+6=14:7
8*6=48:24
Answer 8*6